# Covariant effective spacetimes of spherically symmetric electro-vacuum with a cosmological constant

### v1

@@ IV Covariance of gravity coupled to a general matter field
As did for the EM field, to maintain covariance for the effective metric $g_{\rho\sigma}^{(\mu)}$, the total Hamiltonian constraint needs to share the similar algebraic structure as Eq. (93), and to satisfy the covariance conditions presented in Sec. III.1. Let us assume that this can be implemented by modifying the classical matter Hamiltonian into the effective one $H^{\rm matter}_{\rm eff}$, which incorporates the quantity $\mu$. Since the classical matter Hamiltonian usually depends only on $E^{1}$ and $E^{2}$ as the components of the spatial metric, incorporating $\mu$ into $H^{\rm matter}_{\rm eff}$ does not introduce a dependence on derivatives of $K_{1}$. Thus, the covariance condition (i) is satisfied. Moreover, it can be verified that
@@
Since Eq. (167) remains valid with replacing $H_{\rm eff}^{\rm matter}$ by $H_{\rm eff}^{GR}$ given in Eq. (92). It is concluded that Eq. (167) is satisfied by the total effective Hamiltonian constraint $H_{\rm eff}^{GR}+H_{\rm eff}^{\rm matter}$. Namely, the covariance condition (ii) is also satisfied.
According to the above discussion, to maintain the covariance for the effective metric $g_{\rho\sigma}^{(\mu)}$ in the coupling model with a general matter field whose classical Hamiltonian depends only on $E^{1}$ and $E^{2}$, it is sufficient to modify the classical matter Hamiltonian by incorporating $\mu$ so that the total Hamiltonian constraint retains the similar algebraic structure as Eq. (93).
It is important to note that in the coupling model with matter fields, maintaining covariance also requires the matter fields themselves to be covariant. This needs us to define an effective 4-D matter field that may depend on phase space variables and the Lagrange multiplier, similar to the treatment of the field $A^{(\mu)}_{\rho}$. In general, the precise definition of the effective matter field requires further investigation.
@@ V Summary and discussion
In this paper, we successfully extended the approach introduced in [41, 42] for restoring covariance in the effective quantum BH models to the electro-vacuum case with a cosmological constant. In Sec. II, we first recalled the Hamiltonian framework of gravity coupled to EM field in the spherically symmetric case. Specifically, we introduced the EM diffeomorphism constraint (9) following Refs. [46, 47], such that the total diffeomorphism constraint generates the spatial diffeomorphism transformations for gravity and EM field. Additionally, the EM Gauss constraint is incorporated into the Hamiltonian (7) of EM field, ensuring that the total Hamiltonian and diffeomorphism constraints of the coupling model form a subalgebra (11d)–(11f) as that of the vacuum model in [41, 42]. As a result, the analysis for spacetime covariance in the coupling model will be parallel to that in vacuum case. The covariance condition (26) for gravity and the one (27) for EM field have been examined in the classical theory. By solving the Gauss constraint, we derived the solutions (72) and (76) to the EOM in different gauge fixings, corresponding to the RN-(anti-)de Sitter metric.
@@
In Sec. III, we studied covariance of the effective coupling model. To ensure validity of the constrain algebra (80) and the covariance conditions (i) and (ii), the classical Hamiltonian (7) of EM field has been modified into the effective one (92) by incorporating $\mu$, leading to the covariance of effective metric. To require the effective vector potential satisfying the covariance equation (96), we obtained an explicitly form (99) of $A_{\rho}^{(\mu)}$ by a detailed derivation. Based on the covariance effective coupling model, we further explored its exact solutions. Thanks to the three solutions $M_{\rm eff,I}$, $M_{\rm eff,II}$, and $M_{\rm eff,III}$ to Eq. (85) for building the effective Hamiltonian constraints of gravity in vacuum model, the corresponding exact solutions for the effective coupling model have been founded in Sec. III.3, Sec. III.4, and Sec. III.5. In the solutions corresponding to $M_{\rm eff,I}$, the vector potential $A_{\rho}^{(\mu_{\rm I})}$ remains its classical form (66) in a specific gauge choice, while the effective metric receives the quantum-corrections with the line element (120) in Schwarzschild coordinates and (123) in PG ones. Interestingly, in the solutions corresponding to $M_{\rm eff,II}$ (and $M_{\rm eff,III}$), there exist effective vector potentials that incorporate quantum gravity effects, as shown in Eqs. (151) and (156), due to existence of arbitrary functions $\mathcal{X}(\mu)$ and $\mathcal{Y}(\mu)$. Hence, in these cases, quantum gravity effects could affect EM field, besides changing the structure of spacetime.
@@
Finally, how to extend covariance to a general matter field coupling was discussed. The main result can be summarized as follows: to preserve covariance for the effective metric $g_{\rho\sigma}^{(\mu)}$ in the coupling model with a general matter field whose classical Hamiltonian depends only on $E^{1}$ and $E^{2}$, it is sufficient to modify the classical matter Hamiltonian by incorporating $\mu$, ensuring that the total Hamiltonian constraint maintains the similar algebraic structure as Eq. (93).
Let us conclude this section by comparing our work with the study presented in [51]. The work [51] considered the scalar constraint and the vector constraint as the basis of the constraint algebra. The scalar constraint is a linear combination of the Hamiltonian constraint and the Gauss constraint considered in this paper, while the vector constraint corresponds to a linear combination of the diffeomorphism constraint and the Gauss constraint. The scalar and vector constraints form a subalgebra with the same structure as given in Eqs. (11d)–(11f). However, the vector constraint does not generate diffeomorphism transformation for $A_{\rho}$, but generates the diffeomorphism transformation for the field strength tensor $F_{\mu\nu}$. This is actually easy to explain. The vector constraint generates a diffeomorphism transformation plus a $U(1)$ gauge transformation while $F_{\mu\nu}$ is $U(1)$-gauge invariant. Due to the above facts, $F_{\mu\nu}$ is taken as a basic variable in the work [51] when constructing the effective model. Nevertheless, as demonstrated in this work, this substitution is not necessary. When the diffeomorphism constraint is directly incorporated, it is sufficient to modify the scalar constraint into the Hamiltonian constraint while ensuring that the algebraic structure given in Eqs. (11d)–(11f) is preserved. Since the vector constraint generates the diffeomorphism transformation for $A_{\rho}$, it can be chosen as a fundamental variable in the effective model. Our approach places the classical and effective models on an equal footing. Moreover, it can be easily generalized to non-Abelian Yang-Mills theory, where $F_{\mu\nu}$ is no longer gauge invariant.
@@ Acknowledgements.
This work is supported in part by NSFC Grants No. 12165005 and No. 12275022.

### v3

@@ IV Covariance of gravity coupled to a general matter field
As done for the EM field, to maintain covariance for the effective metric $g_{\rho\sigma}^{(\mu)}$, the total Hamiltonian constraint needs to share the similar algebraic structure as Eq. (80), and to satisfy the covariance conditions presented in Sec. III.1. Let us assume that this can be implemented by modifying the classical matter Hamiltonian into the effective one $H^{\rm matter}_{\rm eff}$, which incorporates the quantity $\mu$. Since the classical matter Hamiltonian usually depends only on $E^{1}$ and $E^{2}$ as the components of the spatial metric and since $\mu$ depends only on $E^{I}$ and $K_{2}$ according to the covariance equation (73), incorporating $\mu$ into $H^{\rm matter}_{\rm eff}$ does not introduce a dependence on derivatives of $K_{1}$. Thus, the covariance condition (i) is satisfied. Moreover, it can be verified that
@@
Since Eq. (166) remains valid with replacing $H_{\rm eff}^{\rm matter}$ by $H_{\rm eff}^{GR}$ given in Eq. (79), it is concluded that Eq. (166) is satisfied by the total effective Hamiltonian constraint $H_{\rm eff}^{GR}+H_{\rm eff}^{\rm matter}$. Namely, the covariance condition (ii) is also satisfied.
According to the above discussion, to maintain the covariance for the effective metric $g_{\rho\sigma}^{(\mu)}$ in the coupling model with a general matter field whose classical Hamiltonian depends only on $E^{1}$ and $E^{2}$, it is sufficient to modify the classical matter Hamiltonian by incorporating $\mu$ and requiring that the total Hamiltonian constraint retains the similar algebraic structure as Eq. (80). As discussed in Appendix B, Eqs. (80d) and (80e) follow directly from the fact that the diffeomorphism constraint generates spatial diffeomorphism transformations. Therefore, as long as the classical form of the diffeomorphism constraint is preserved, these two equations hold automatically. Consequently, our attention can be focused on the final equation, Eq. (80f). Based on the derivations in Appendix B, the cross term
$\{H_{\rm eff}^{\rm matter}[N_{1}]+H^{GR}_{\rm eff}[N_{1}],H_{\rm eff}^{\rm matter}[N_{2}]+H^{GR}_{\rm eff}[N_{2}]\}$
vanishes since $H_{\rm eff}^{\rm matter}$ depends on $E^{I}$ and $\mu$. In addition, the bracket $\{H^{\mathrm{GR}}_{\mathrm{eff}}[N_{1}],H^{\mathrm{GR}}_{\mathrm{eff}}[N_{2}]\}$ already takes the correct form as given in Eq. (80f). Therefore, to ensure that the full constraint algebra retains the correct structure, it is sufficient to require that the modified matter Hamiltonian have the correct Poisson bracket structure with itself.
It is important to note that in the coupling model with matter fields, maintaining covariance also requires the matter fields themselves to be covariant. This needs us to define an effective 4D matter field that may depend on phase space variables and the Lagrange multiplier, similar to the treatment of the field $A^{(\mu)}_{\rho}$. In general, the precise definition of the effective matter field requires further investigation.
@@ V Summary and discussion
In this paper, we successfully extended the approach introduced in [42, 43] for restoring covariance in the effective quantum BH models to the electrovacuum case with a cosmological constant. In Sec. II, we first recalled the Hamiltonian framework of gravity coupled to the EM field in the spherically symmetric case. Specifically, we introduced the EM diffeomorphism constraint (9) following Refs. [49, 50], such that the total diffeomorphism constraint generates the spatial diffeomorphism transformations for gravity and the EM field. Additionally, the EM Gauss constraint is incorporated into the Hamiltonian (7) of the EM field, ensuring that the total Hamiltonian and diffeomorphism constraints of the coupling model form a subalgebra (11d)–(11f) as that of the vacuum model in [42, 43]. As a result, the analysis for spacetime covariance in the coupling model will be parallel to that in the vacuum case. The covariance condition (24) for gravity and the one (25) for the EM field have been examined in the classical theory. By solving the Gauss constraint, we derived the solutions (59) and (63) to the EOM in different gauge fixings, corresponding to the RN–(anti–)de Sitter metric.
@@
In Sec. III, we studied covariance of the effective coupling model. To ensure validity of the constraint algebra (67) and the covariance conditions (i) and (ii), the classical Hamiltonian (7) of the EM field has been modified into the effective one (79) by incorporating $\mu$, leading to the covariance of the effective metric. To require the effective vector potential satisfying the covariance equation (83), we obtained an explicit form (86) of $A_{\rho}^{(\mu)}$ by a detailed derivation. Based on the covariance effective coupling model, we further explored its exact solutions. Thanks to the three solutions $M_{\rm eff,I}$, $M_{\rm eff,II}$, and $M_{\rm eff,III}$ to Eq. (73) for building the effective Hamiltonian constraints of gravity in the vacuum model, the corresponding exact solutions for the effective coupling model have been founded in Sec. III.3, Sec. III.4, and Sec. III.5. In the solutions corresponding to $M_{\rm eff,I}$, the vector potential $A_{\rho}^{(\mu_{\rm I})}$ remains its classical form (53) in a specific gauge choice, while the effective metric receives the quantum corrections with the line element (107) in Schwarzschild coordinates and (110) in PG ones. Interestingly, in the solutions corresponding to $M_{\rm eff,II}$ (and $M_{\rm eff,III}$), there exist effective vector potentials that incorporate quantum gravity effects, as shown in Eqs. (138) and (143), due to the existence of arbitrary functions $\mathcal{X}(\mu)$ and $\mathcal{Y}(\mu)$. Hence, in these cases, quantum gravity effects could affect the EM field, besides changing the structure of spacetime. Moreover, the influence of quantum gravity effects on the spacetime structure has been preliminarily explored in Secs. III.3.1, III.4.4, and III.5.1, focusing on the case with a vanishing cosmological constant, a small charge, and a large black hole mass. Although only a preliminary investigation, this analysis raises several open questions, for example, how to construct Penrose diagrams for spacetimes with multiple horizons, and what the implications are of the singularity resolution in ${\rm d}s_{\rm eff,II}^{2}$ and ${\rm d}s_{\rm eff,III}^{2}$. A more comprehensive study, including an investigation of the spacetime structure in other regimes with a nonvanishing cosmological constant and a large charge, will be valuable, and we leave these questions for future work.
@@
Finally, how to extend covariance to a general matter field coupling was discussed in Sec. IV. The main result can be summarized as follows: to preserve covariance for the effective metric $g_{\rho\sigma}^{(\mu)}$ in the coupling model with a general matter field whose classical Hamiltonian depends only on $E^{1}$ and $E^{2}$, it is sufficient to modify the classical matter Hamiltonian by incorporating $\mu$, ensuring that the total Hamiltonian constraint maintains the similar algebraic structure as Eq. (80).
@@ Acknowledgements.
This work is supported in part by NSFC Grants No. 12165005, No. 12275022, and “the Fundamental Research Funds for the Central Universities”.
Appendix A Covariance of the classical theory
Observing that the change $\delta g_{\rho\sigma}$ of $g_{\rho\sigma}$ given by Eq. (20) includes $\delta E^{I}$, $\delta N$, and $\delta N^{x}$, let us first calculate $\delta E^{I}$ by definition (27). As discussed in [42, 43], it is sufficient to consider the case where $\alpha$, $\beta^{x}$, and $\lambda$ are phase space independent. Then, for $\alpha$, $\beta^{x}$, and $\lambda$ being phase space dependent, all of the results will still apply, but only on-shell, i.e., when the constraints are satisfied. We have
$\delta E^{1}$ $=\mathcal{L}_{\alpha\mathfrak{N}+\beta}E^{1},$ (169)
To calculate $\delta N$ and $\delta N^{x}$, we consider that the gauge transformation (26) will map a solution $K_{I}(t,x)$, $E^{I}(t,x)$ to a new set of solutions
As discussed in Sec. II.2, the transformed solutions must satisfy Hamilton’s equations with respect to the modified lapse function $\tilde{N}=N+\epsilon\,\delta N$ and shift vector $\tilde{N}^{x}=N^{x}+\epsilon\,\delta N^{x}$. It is worth noting that $\delta N$ and $\delta N^{x}$ were derived in the vacuum case in [42, 43]. These results, as presented by the first two equations in Eq. (29), can be directly adopted here, as they rely on two key ingredients that remain intact in the current nonvacuum model: (a) the fact that the diffeomorphism constraint generates spatial diffeomorphism transformations, which results in EOM for $E^{I}$ and $K_{I}$ taking the same forms as Eq. (23); and (b) the structure of constraint algebra, which is the same as Eqs. (11d)–(11f).
Similar to getting Eq. (169), we have
$\tilde{\Gamma}$= $\Gamma+\epsilon\delta\Gamma+O(\epsilon^{2}),$ (172)
They need to satisfy Hamilton’s equations with respect to a new lapse function $\tilde{N}=N+\epsilon\delta N$, a new shift vector $\tilde{N}^{x}=N^{x}+\epsilon\delta N^{x}$, and a new Lagrangian multiplier $\tilde{\Psi}=\Psi+\epsilon\delta\Psi$, where $\delta N$ and $\delta N^{x}$ have been given by Eq. (29).
Substituting $\tilde{\Gamma}$ and $\tilde{P}_{\Gamma}$ given by Eq. (172) into Eq. (23) and comparing the first order results in $\epsilon$ on both sides, we get
$+\Big\{\left\{X,H[N]+\mathcal{G}[\Psi]\right\},H[\alpha N]+H_{x}[\beta^{x}]+\mathcal{G}[\alpha\Psi+\lambda]\Big\},$
Then, we compute the second term on the left-hand side of Eq. (173) and get
$\mathcal{L}_{\mathfrak{N}}\delta X$ (175)
$+\{X,H[\mathcal{L}_{\mathfrak{N}}(\alpha N)]+H_{x}[\mathcal{L}_{\mathfrak{N}}\beta^{x}]+\mathcal{G}[\mathcal{L}_{\mathfrak{N}}(\alpha\Psi+\lambda)]\}.$
Note that in the last step in Eq. (175), the first term contains the Lie derivatives for the phase-space-dependent parts in $\delta X$, and the second term includes the Lie derivative for the phase-space-independent variables. Applying the Jacobi identity for Poisson brackets to Eq. (175), we get
$\mathcal{L}_{\mathfrak{N}}\delta X=$ $\Big\{\left\{X,H[N]+\mathcal{G}[\Psi]\right\},H[\alpha N]+H_{x}[\beta^{x}]+\mathcal{G}[\alpha\Psi+\lambda]\Big\}$ (176)
$+\Big\{X,\left\{H[\alpha N]+H_{x}[\beta^{x}]+\mathcal{G}[\alpha\Psi+\lambda],H[N]+\mathcal{G}[\Psi]\right\}\Big\}$
$+\{X,H[\mathcal{L}_{\mathfrak{N}}(\alpha N)]+H_{x}[\mathcal{L}_{\mathfrak{N}}\beta^{x}]+\mathcal{G}[\mathcal{L}_{\mathfrak{N}}(\alpha\Psi+\lambda)]\}.$
The second term on the right-hand side of Eq. (176) contains a Poisson bracket between constraints, and thus it can be evaluated by applying the constraint algebra (80a)–(80f). The result reads
$\mathcal{L}_{\mathfrak{N}}\delta X=$ $\Big\{\left\{X,H[N]+\mathcal{G}[\Psi]\right\},H[\alpha N]+H_{x}[\beta^{x}]+\mathcal{G}[\alpha\Psi+\lambda]\Big\}$ (177)
$+\left\{X,\mathcal{G}\left[N\sqrt{S}(-\Psi\partial_{x}\alpha-\partial_{x}\lambda)+\beta^{x}\partial_{x}\Psi\right]\right\}$
$+\Big\{X,H[\mathcal{L}_{\mathfrak{N}}(\alpha N)]+H_{x}[\mathcal{L}_{\mathfrak{N}}\beta^{x}]+\mathcal{G}[\mathcal{L}_{\mathfrak{N}}(\alpha\Psi+\lambda)]\Big\}.$
Putting Eqs. (174) and (177) into Eq. (173) yields
$+\Big\{X,H[\beta^{x}\partial_{x}N+\mathcal{L}_{\mathfrak{N}}(\alpha N)-\delta N]\Big\}$
$+\left\{X,\mathcal{G}\left[\xi\right]\right\}=0,$
where the coefficient $\xi$ is
$\xi=-N\sqrt{S}(\Psi\partial_{x}\alpha+\partial_{x}\lambda)+\beta^{x}\partial_{x}\Psi+\mathcal{L}_{\mathfrak{N}}(\alpha\Psi+\lambda)-\delta\Psi,$ (179)
Note that the first two terms in Eq. (178) vanish due to Eq. (29). Thus, Eq. (178) gives
Appendix B Detained calculations on Eq. (80)
Since the Gauss constraint $\mathcal{G}$ contains only $P_{\Gamma}$, Eqs. (80a), (80b), and (80c) can easily be checked. For Eqs. (80d) and (80e), the results are just a direct consequence that $H_{x}$ generates spatial diffeomorphism transformation. Specifically, using the expression of $H_{x}=H_{x}^{GR}+H_{x}^{EM}$ given in Eqs. (8) and (9), we have
$\{O,H_{x}[N^{x}]\}=$ $N^{x}\partial_{x}O=\mathcal{L}_{N^{x}\partial_{x}}O,\forall O=K_{2},E^{1},P_{\Gamma},$ (181)
$\{\tilde{O},H_{x}[N^{x}]\}=$ $\partial_{x}(N^{x}\tilde{O})=\mathcal{L}_{N^{x}\partial_{x}}\tilde{O},\forall\tilde{O}=E^{2},K_{1},\Gamma,\partial_{x}P_{\Gamma}.$
where we should note that $\tilde{O}$ are scalar density with weight 1. As a consequence of these equations, we have
$\{s_{a},H_{x}[N^{x}]\}=N^{x}\partial_{x}s_{a}=\mathcal{L}_{N^{x}\partial_{x}}s_{a},\forall a=3,4,5,$ (182)
where the second equality results from that $s_{a}$ are scalars. According to Eqs. (181) and (182), for any phase space function $F$ depending on $v^{I}\in\{P_{\Gamma},\Gamma,E^{2},K_{1},s_{1}\equiv E^{1},s_{2}\equiv K_{2},s_{3},s_{4},s_{5}\}$, we get
$\{F,H_{x}[N^{x}]\}=\frac{\partial F}{\partial v^{I}}\{v^{I},H_{x}[N^{x}]\}=\mathcal{L}_{N^{x}\partial_{x}}F.$ (183)
Clearly, $H_{x}$ and $H_{\rm eff}$ are such a functions. Thus,we get
$\{H_{x}[N_{1}^{x}],H_{x}[N_{2}^{x}]\}=$ $\int{\rm d}xN_{1}^{x}\mathcal{L}_{N_{2}^{x}\partial_{x}}H_{x}$ (184)
$=$ $\int{\rm d}xN_{1}^{x}\left(N_{2}^{x}\partial_{x}H_{x}+2H_{x}\partial_{x}N_{2}^{x}\right),$
$\{H_{x}[N^{x}],H_{\rm eff}[N]\}=$ $-\int{\rm d}xN\mathcal{L}_{N^{x}\partial_{x}}H_{\rm eff}$
$=$ $-\int{\rm d}xN\partial_{x}\left(N^{x}H_{\rm eff}\right),$
which is the same as Eqs. (80d) and (80e) up to a boundary term. The only remaining equation that requires further effort is Eq. (80f).
For Eq. (80f), due to
$\{O[N_{1}],O[N_{2}]\}=H_{x}[\mu S(N_{1}\partial_{x}N_{2}-N_{2}\partial_{x}N_{1})],\,\forall O=H_{\rm eff}^{GR},\,H_{\rm eff}^{EM},$
we only need to show the cross term in $\{H_{\rm eff}[N_{1}],H_{\rm eff}[N_{2}]\}$ vanishes. Namely, we need to show
$\{H_{\rm eff}^{GR}[N_{1}],H_{\rm eff}^{EM}[N_{2}]\}+\{H_{\rm eff}^{EM}[N_{1}],H_{\rm eff}^{GR}[N_{2}]\}=0.$ (185)
Observing that $H_{\rm eff}^{GR}$ does not depend on $\Gamma$ and $P_{\Gamma}$, we have, according to Eq. (79) which indicates that $H_{\rm eff}^{EM}$ is a function of $\mu$, $E^{I}$, $\Gamma$, $P_{\Gamma}$, and $\partial_{x}P_{\Gamma}$,
$\{H_{\rm eff}^{GR}[N_{1}],H_{\rm eff}^{EM}[N_{2}]\}+\{H_{\rm eff}^{EM}[N_{1}],H_{\rm eff}^{GR}[N_{2}]\}$ (186)
$=$ $\sum_{I=1,2}\int{\rm d}xN_{2}(x)\frac{\partial H_{\rm eff}^{EM}}{\partial E^{I}}\left\{H_{\rm eff}^{GR}[N_{1}],E^{I}\right\}$
$+\int{\rm d}xN_{2}(x)\frac{\partial H_{\rm eff}^{EM}}{\partial\mu}\left\{H_{\rm eff}^{GR}[N_{1}],\mu\right\}$
$-N^{1}\leftrightarrow N^{2}.$
The notation $N_{1}\leftrightarrow N_{2}$ indicates that the final line is obtained by exchanging $N_{1}$ and $N_{2}$ in the preceding terms. Since $H_{\rm eff}^{GR}$ does not depend on the derivatives of $K_{I}$ for $I=1,2$, we have, for the first term in the right hand side of Eq. (186),
$N_{2}(x)\frac{\partial H_{\rm eff}^{EM}}{\partial E^{I}}\left\{H_{\rm eff}^{GR}[N_{1}],E^{I}\right\}\propto N_{1}(x)N_{2}(x),$ (187)
which will be exactly canceled by its counterpart in the antisymmetric combination appearing in the expression with $N_{1}\leftrightarrow N_{2}$. This result leads us to simplify Eq. (186) to be
$\{H_{\rm eff}^{GR}[N_{1}],H_{\rm eff}^{EM}[N_{2}]\}+\{H_{\rm eff}^{EM}[N_{1}],H_{\rm eff}^{GR}[N_{2}]\}$ (188)
$=$ $\int{\rm d}xN_{2}(x)\frac{\partial H_{\rm eff}^{EM}}{\partial\mu}\left\{H_{\rm eff}^{GR}[N_{1}],\mu\right\}-N^{1}\leftrightarrow N^{2}$
$=$ $\left(\{H_{\rm eff}^{GR}[N_{1}],A[\mu]\}-N_{1}\leftrightarrow A\right)\Big|_{A=N_{2}(x)\frac{\partial H_{\rm eff}^{EM}}{\partial\mu}},$
where $A$ denotes a phase-space-independent scalar density. We thus need to show $\{H_{\rm eff}^{GR}[N_{1}],A[\mu]\}-N_{1}\leftrightarrow A$ vanishes. Noting that $\mu$ depends on $s_{1}$, $s_{2}$, and $s_{4}$, we have
$\{H_{\rm eff}^{GR}[N_{1}],A[\mu]\}-N_{1}\leftrightarrow A$ (189)
$=$ $-\int{\rm d}x{\rm d}yA(x)N_{1}(y)(\partial_{s_{2}}\mu)(x)\left(\partial_{s_{5}}H_{\rm eff}^{GR}\right)(y)\{s_{2}(x),s_{5}(y)\}$
$+\int{\rm d}x{\rm d}yN_{1}(x)A(y)\left(\partial_{s_{3}}H_{\rm eff}^{GR}\right)(x)(\partial_{s_{4}}\mu)(y)\{s_{3}(x),s_{4}(y)\}$
$-N_{1}\leftrightarrow A,$
where we have omitted terms involving Poisson brackets of the form $\{s_{a}(y),s_{b}(x)\}\propto\delta(x,y)$, as these lead to contributions proportional to $A(x)N_{1}(x)$ and thus are canceled under the antisymmetrization $N_{1}\leftrightarrow A$. A direct calculation yields
$\{s_{2}(x),s_{5}(y)\}$ $=\left[3G\frac{[\partial_{y}E^{1}(y)][\partial_{y}E^{2}(y)]}{[E^{2}(y)]^{4}}-2G\frac{\partial_{y}^{2}E^{1}(y)}{[E^{2}(y)]^{3}}\right]\delta(x,y)$ (190)
$\quad-G\frac{\partial_{y}E^{1}(y)}{[E^{2}(y)]^{3}}\partial_{y}\delta(x,y),$
and
$\{s_{3}(x),s_{4}(y)\}$ $=\frac{2G}{E^{2}(x)E^{2}(y)}\partial_{y}\delta(x,y).$ (191)
Substituting the results into (189), we get
$\{H_{\rm eff}^{GR}[N_{1}],A[\mu]\}-N_{1}\leftrightarrow A$ (192)
$=$ $-G\int{\rm d}x\left(A\partial_{x}N_{1}-N_{1}\partial_{x}A\right)\frac{\partial_{x}E^{1}(\partial_{s_{2}}\mu)\partial_{s_{5}}H_{\rm eff}^{GR}}{[E^{2}]^{3}}$
$-G\int{\rm d}x{\rm d}y\left(N_{1}\partial_{x}A-A\partial_{x}N_{1}\right)\frac{2\left(\partial_{s_{3}}H_{\rm eff}^{GR}\right)(\partial_{s_{4}}\mu)}{[E^{2}]^{2}},$
According to Eqs. (72) and (73b), we have
$\frac{\partial_{x}E^{1}(\partial_{s_{2}}\mu)\partial_{s_{5}}H_{\rm eff}^{GR}}{[E^{2}]^{3}}=\frac{-2(\partial_{s_{2}}\mu)(\partial_{s_{4}}M_{\rm eff})}{E^{2}}=\frac{2\left(\partial_{s_{3}}H_{\rm eff}^{GR}\right)(\partial_{s_{4}}\mu)}{[E^{2}]^{2}}.$ (193)
Inserting this result into Eq. (192) leads to that $\{H_{\rm eff}^{GR}[N_{1}],A[\mu]\}-N_{1}\leftrightarrow A$ vanishes, as what we wanted.